A home gym is listed for $786.20 less 36%, 10%, and 2%. (a) What is the net price? (b) What is the total amount of discount allowed? (c) What is the exact single rate of discount that was allowed?


1
Expert's answer
2021-01-13T18:42:41-0500

a) Let the net price be represented by "N," the original list price by "L," the first rate of discount by "d_1," the second rate of discount by "d_2," the third rate of discount by "d_3"

"N=L(1-d_1)(1-d_2)(1-d_3)"

"N=\$786.20(1-0.36)(1-0.10)(1-0.02)""=\$443.79"

The net price is "\$443.79."


b)


"Amount\ of\ discount=L-N"

"=\$786.20-\$443.79=\$342.41"

The total amount of discount allowed is "\$342.41."


c)


"\text{Single equivalent rate of discount}"

"=1-(1-d_1)(1-d_2)(1-d_3)"

"=1-(1-0.36)(1-0.10)(1-0.02)"

"=0.43552"

The exact single rate of discount that was allowed is "43.552\%."